# Phase Retrieval by Quaternionic Reweighted Amplitude Flow on Image Reconstruction

### v1

@@ 4.3 Adapted Algorithm: QAdRAF
0:  Same as in QRAF. Moreover, set the coefficient $\mu\in\left(0,1\right)$ and parameter $\epsilon$;
0:  $\boldsymbol{z}_{T}$;
@@
4:     Let $S_{0}=0$ and $S_{i+1}=\mu S_{i}+\left(1-\mu\right)\left|{\nabla f(\boldsymbol{z}_{i})}\right|^{2}$, then we compute
@@ 5 Quaternionic Perturbed Amplitude Flow (QPAF)
$\min_{\boldsymbol{z}}f_{\varepsilon}(\boldsymbol{z})=\min_{\boldsymbol{z}}\frac{1}{n}\sum_{j=1}^{n}\left(\sqrt{|\alpha_{j}^{*}\boldsymbol{z}|^{2}+\varepsilon^{2}}-\sqrt{b_{j}^{2}+\varepsilon_{j}^{2}}\right)^{2},$ (5.1)
where $\boldsymbol{\varepsilon}=[\varepsilon_{1},\cdots,\varepsilon_{n}]\in\mathbb{R}^{n}$ is a vector with prescribed values, subject to the condition that $\varepsilon_{j}\neq 0$ for all $b_{j}\neq 0$. When $\varepsilon_{j}=0$ for all $j=1,\cdots,n$, this model reduces to the amplitude-based model (1.2), and is therefore referred to as the perturbed amplitude-based model. Based on this formulation, a new non-convex algorithm, called the perturbed Amplitude Flow (PAF) is developed. The PAF algorithm can recover the target signal under $\mathcal{O}(d)$ Gaussian random measurements. Starting from a well-designed initial point, PAF converges at a linear rate for both real and complex signals, without requiring truncation or reweighted procedures, while maintaining strong empirical performance.
@@
3:  The initialization $\boldsymbol{z}_{0}=\lambda_{0}\cdot\boldsymbol{\nu}_{in}$..
@@ 6 Algorithm regarding to RGB picture
In quaternion image processing, the color channels (Red, Green, and Blue channels) are usually represented by the three imaginary components of quaternions, making the desired signal taking pure quaternions. Many studies (see e.g., [35]) opt to eliminate the real part of the quaternion signals after reconstruction. However, this approach generally performes poorly in the context of phase retrieval for pure quaternion signals. A more effective approach is to incorporate the pur quaternion assumption directly into the recovery process, as done in [9]. In this section, we brief review of the Phase Factor Estimate (QPFE) technique used in [9], which will be applied to refine our algorithms introduced in previous sections and then applied to phase retrieval of natural images.
@@
5:     Compute the left product of $\boldsymbol{z}$ by $\bar{q}$, i.e.
$\boldsymbol{p}=\boldsymbol{z}\cdot\bar{q}.$
@@
9:  return $\boldsymbol{z}_{T}$.
Similar to [9], we use $T_{p}$ to be the parameter that controls the transform of full quaternion to pure quaternion signals. Unlike [9] which utilizes inner and outer loop for the result, we only perform QPFE once after $T_{p}$ iteration periodically. In our algorithm, the total iteration number is the given parameter $T$ compare to the actual iteration number $T\times T_{p}$ in [9].
It is easily to see that the parameter $T_{p}$ makes difference when it changes, especially when $T_{p}=1$ which means operate QPFE in every iteration and $T_{p}=T$ which means operate QPFE at the final step of the full quaternion algorithm. Although there are difference among them, the actual computation shows no essential distinction. Since our work considers mainly on different algorithms rather than different aspects of a single algorithm, we prefer to save this part of work in the future.
@@ 7 Numerical Experiments
In this section, we show the numerical efficiency of QRAF and its improvements by comparing their the performance with other competitive methods. Specifically, Section 7.1 is for synthetic data, while Section 7.2 focuses on color image processing. Additionally, Section 7.3 provides a comparative analysis of conventional and quaternion-based algorithms.
@@
Algorithm Parameters
QWF $\eta=\frac{0.2n}{\sum_{k=1}^{n}y_{k}}$;
QRWF $\alpha_{\ell}=1$, $\alpha_{u}=5$ and $\mu=0.8$;
QPAF $\gamma=1/2$, $\eta=2.5$ and $\sigma=2$;
QTAF $\gamma=0.8$, $\eta=1.2$ and $\rho=\frac{1}{6}$;
QRAF $\left|{\mathcal{S}}\right|=\lfloor 3n/13\rfloor$, $\beta=5$, $\gamma=0.5$ and $\mu=6$;
QARAF $\eta=6$ and $\mu=0.8$;
QAdRAF $\alpha=0.009$, $\mu=1$, $\epsilon=10^{-6}$;
QIRAF $\eta=6$ and batch size $\left|{\Gamma}\right|=2^{k}$, $k=\frac{\log\left(\frac{n}{4}-1\right)}{\log 2}$.
Throughout our work, the parameters of the tested algorithms are listed in Table 1. For the existing quaternionic algorithm QWF, we use the given parameters there. For the algorithms that transferred from the real or complex space, we compare the original parameters with a huge a mount of tested parameters to obtain a suitable set of parameters.
@@ 7.1 Synthetic Data
i. Convergence performance   Firstly, we examine the convergence of each algorithm, as shown in Fig. 2 and Fig. 2. The relative error, defined as $\text{dist}(x,z)/\left\|{x}\right\|$ and $\text{dist}_{p}(x,z)/\left\|{x}\right\|$ w.r.t full and pure quaternion algorithms respectively, serves as the convergence criterion. It is seen that all algorithms are convergent albeit at different rates. The convergence rate is a crucial aspect and is influenced not only by the algorithm’s gradient but also by other factors, such as the vector dimension $d$, the ratio $n/d$, and the algorithm’s initialization.
@@
In this sample, we set $n/d=9$ to simply investigate its convergence. It is easy to see that in general all algorithms converge to the same level of relative error, i.e. $\ln(10^{-35})$, but there are some details that must be noted.
For the full quaternion algorithms in Fig. 2, QARAF leads the competition of convergence with the iteration number less than $400$. QRAF and its variants converge evidently faster than other algorithms. The QWF also converges but with a lower speed that is not fully shown in the figure. We should note that the curves of QRAF and QIRAF almost cover each other, showing the same ability of convergence. This is actually important since as the variants of QRAF, QIRAF has the advantage of computation speed which we will discuss later in details.
@@
ii. Success rate  Next, we report the success rates of the QRAF, QRWF, QTAF, QWF and QIRWF algorithms under different sample size. Specifically, for a vector $\boldsymbol{x}\in\mathbb{H}^{d}$ with $d=100$, we conduct $100$ trials for each value of $n/d$, ranging from $3$ to $13$ in increments of $0.5$. Taking into account $85<n/d<$ 8 is critical, the step size was reduced to $0.2$ to provide clear insights in the corresponding figure. Each algorithm was run for 1500 iterations per trial. Similar as [9, 17], a trial was deemed successful if it achieved $\text{dist}(\boldsymbol{x},\boldsymbol{z})<10^{-5}$ within the 1500 iterations. The parameters used for each algorithm are listed in the Table 1, and the success rate regarding to $n/d$ is in Fig. 4 and Fig. 4.
Figure 3: Success rate of different quaternion algorithms
Figure 4: Success rate of QRAF and its variants
In order to show results in detail, we split the algorithms in two sets, i.e. in Fig. 4 and Fig. 4. For the comparison of Wirtinger flow (QWF, QRWF) and Amplitude flow (QTAF, QRAF, QPAF), it can be easily seen that QRAF performs an overall advantage in success rate than other algorithms. Meanwhile, QPAF performs also well in this index with a success rate equals to $1$ when $n/d$ equals to $8$. This value is larger than $10$ in algorithms such as QWF and QTAF. Since $n/d$ rate is decisive in computation cost, for the same success rate, a lower $n/d$ value means less computation, higher speed and less computer memory requirement. Thus in this point of view, QRAF outperform other algorithms in this set.
However, in the set of QRAF and its variants in Fig. 4, there is another story. The most attention-tracking algorithm is QARAF which reduces the $n/d$ rate to $7$ to get the success rate of $1$and outperformed other algorithms. This algorithm enjoys an accelerated steepest decent method as introduced in subsection 4.2. The QAdRAF with an adapted gradient decent method also performed better than QRAF, thus better than other algorithms in the previous set in Fig. 4. Critically, the QIRAF seems no difference with QRAF which is the important. With the mini-batch strategy, QIRAF decrease huge amount of computation but loose almost nothing in success rate. This enables QIRAF performed as fast (see Table 2) as QARAF which plays an important role in in practical.
It worth mention two things in this index (success rate). Firstly, since we define a success as the algorithm achieves $\text{dist}(\boldsymbol{x},\boldsymbol{z})<10^{-5}$ within 1500 iterations, for a single trial, one can not tell whether the algorithm achieves this threshold after 100 iterations or 1400 iterations. Secondly, there are points of inflection on the curve, showing the oscillation of success rate. The main reason is that the measurement matrix is drawn with randomness and $n/d$ rates are very close (merely 0.2) among the points of inflection. It is easy to understand that if we increase the step size in $n/d$ rate or run more trials, the inflection points will disappear. But both remarks show that the index of success rate is more of a general comparison of algorithms rather than detailed evaluation. Thus the step size and the trial amount that we have chosen is enough for the study.
@@
From above figures, we can easily seen that the signal dimension also affect tested algorithms on convergence at some level. For instance, when $d=30$ and $n/d=6$, the success rate of QIRAF is $0.83$ while when $d=300$ the success rate is 00 with the same $n/d$. Another example is when $d=60$ and $n/d=6.5$ the success rate is $1$, but when $d=300$ and $n/d=6.5$, the success rate is $0.33$. (The original data is uploaded and can be verified by any means.)
In another word, as $d$ increases, the ratio $n/d$ must also be larger to achieve a high success rate. It is worth noting that, compared to QRAF algorithm, the QIRAF is more sensible to the size of vector $\boldsymbol{x}\in\mathbb{H}^{d}$. And the QARAF is the most stable one, comparing to its counterparts.
@@ 7.2 Experiment: Color Images
In this section, we compare different quaternionic algorithms in processing real images. The images used for testing are sourced from the Kodak dataset [^1] https://www.kaggle.com/datasets/sherylmehta/kodak-dataset, i.e. 4.1.04 ($256\times 256$), 4.1.05 ($512\times 512$) and 4.2.03 ($256\times 256$).
Ideally, the entire image would be considered as the signal for reconstruction. However, due to computational limitations present not only in quaternionic algorithms but also in real and complex algorithms, we adopt the conventional approach of segmenting each image into smaller blocks and treating each block as a separate signal $\boldsymbol{x}$ to be reconstructed.
Specifically, in the first experiment, i.e. Fig. 6, we test the picture 4.1.04 in Kodak dataset as in [9]. The picture is $256\times 256$, we split this picture into $32\times 32$ pieces, then each piece is a $8\times 8$ block containing $64$ pixel which is the signal $\boldsymbol{x}$ and its size $d=64$. In the second experiment, i.e. Fig. 7, the figure 4.1.05 of the size $512\times 512$ is also tested. The figure is divided also into $32\times 32$ pieces but with each block of $16\times 16$ pixels. Therefore, in this experiment, the size of the signal $\boldsymbol{x}$ is $d=256$.
For these experiments, we set the $n/d$ ratio to $9$ and fixed the total number of iterations to $T=300$ for each algorithm. The outcomes are presented in Fig. 6.
@@
The results in Fig. 6 clearly highlight the performance of each algorithm. In comparison with the original image, only the reconstruction by PQRAF (see Fig. 6f) contains no defective blocks, indicating the robustness of this algorithm. PQPAF and PQRWF also perform well, with only a few defective blocks, reflecting the varying strengths of WF and AF. Conversely, PQWF and PQTAF exhibit relatively weaker performance, illustrating that different WF and AF algorithms have trade-offs in terms of convergence speed and overall effectiveness.
To quantify these observations, we utilize two evaluation metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). PSNR measures the similarity between the original and reconstructed images, with higher values indicating closer similarity. SSIM assesses structural similarity, ranging from 00 to $1$, where a score of $1$ denotes perfect similarity.
The indices reveal that PQRAF significantly outperforms other algorithms, achieving the highest PSNR score of $96$ in this experiment. Additionally, both PQRAF and PQPAF attain the maximum SSIM score of $1$, leading the competition in this aspect. These indices corroborate the earlier qualitative observations with a more precise, quantitative evaluation.
@@ 7.3 Comparison of conventional and quaternionic algorithms
For a pure quaternion signal $\boldsymbol{p}\in\mathbb{H}^{d}_{p}$, the RAF of a monochromatic model separately reconstruct $\mathcal{P}^{i}(\boldsymbol{p})$, $\mathcal{P}^{j}(\boldsymbol{p})$ and $\mathcal{P}^{k}(\boldsymbol{p})$, which are three real signal in $\mathbb{R}^{d}$. We denote by $\hat{\boldsymbol{z}}^{i}$, $\hat{\boldsymbol{z}}^{j}$ and $\hat{\boldsymbol{z}}^{k}$ the reconstructed signal corresponding to each channel. To keep the metric consistent, the reconstruction error for real RAF of a quaternion signal $\boldsymbol{p}$ is measured as $\left(\sum_{h=i,j,k}{\rm dist}_{p}\left(\mathcal{P}^{h}(\boldsymbol{p}),\hat{\boldsymbol{z}}^{h}\right)^{2}\right)^{1/2}$.
On the other hand, the concatenation model of real RAF for a pure quaternion signal $\boldsymbol{p}\in\mathbb{H}^{d}_{p}$ also split the signal as $\mathcal{P}^{i}(\boldsymbol{p})$, $\mathcal{P}^{j}(\boldsymbol{p})$ and $\mathcal{P}^{k}(\boldsymbol{p})$, and concatenate the three real vector into a new signal, i.e. $\hat{\boldsymbol{p}}=\left[\left(\mathcal{P}^{i}(\boldsymbol{p})\right)^{T},\left(\mathcal{P}^{i}(\boldsymbol{p})\right)^{T},\left(\mathcal{P}^{i}(\boldsymbol{p})\right)^{T}\right]^{T}\in\mathbb{R}^{3d}$. We denote by $\hat{\boldsymbol{z}}\in\mathbb{R}^{3d}$ the reconstructed signal, then the measurement for reconstruction error of this algorithm is defined as ${\rm dist}_{p}\left(\hat{\boldsymbol{p}},\hat{\boldsymbol{z}}\right)$.
@@
However, one thing that cannot be neglected is that real RAF algorithms are evidently faster than quaternionic ones in reconstructing a image as we already mentioned. By simply increasing the $n/d$ rate without any time consumption piles up, real algorithms can surpass the quaternionic ones in convergence as well. However, such comparisons would no longer adhere to the same standard, and thus fall outside the scope of this study.
8 Conclusion
In this paper, we address the quaternionic phase retrieval problem by systematically developing the Quaternionic Reweighted Amplitude Flow (QRAF) algorithm based on an amplitude-based model. The QRAF is further improved by three distinct variants. Numerical results, using both synthetic data and real images, show that our algorithms substantially improve recovery performance and computational efficiency compared to state-of-the-art approaches.
One limitation of QRAF is the absence of a proof for the LCR condition as present in Problem 3.5. This challenge is not unique to the quaternionic domain but also persists in the complex domain. We believe that once the LCR condition is established for complex signals, the proof can be extended to the quaternionic setting without substantial difficulty. Alternatively, to address the difficulty, we considered a perturbed amplitude-based model and developed a non-convex QPAF algorithm. QPAF exhibits comparable performance to QRAF while has theoretical guarantee of linear convergence, which can be extended from proof for complex case.
Furthermore, it is important to note that while quaternionic algorithms offer clear advantages in terms of convergence steps and success rate, their drawback in terms of computational time compared to real-valued algorithms cannot be overlooked. The current implementation of quaternionic algorithms relies on matrix computation techniques. Although the execution speed of these implementations may not be as competitive as real-valued algorithms—largely due to optimizations like MATLAB’s highly accelerated matrix operations—this does not impact the computational complexity or the fundamental priorities of the algorithms under these conditions.

### v2

@@ 4.3 Adapted Algorithm: QAdRAF
0:  Same as in QRAF. Moreover, set the coefficient $\mu\in\left(0,1\right)$, parameters $\epsilon$ and $\alpha$;
@@
4:     Let $S_{0}=0$ and $S_{i+1}=\mu S_{i}+\left(1-\mu\right)\|\nabla f(\boldsymbol{z}_{i})\|^{2}$, then we compute
@@ 5 Quaternionic Perturbed Amplitude Flow (QPAF)
$\min_{\boldsymbol{z}}f_{\varepsilon}(\boldsymbol{z})=\min_{\boldsymbol{z}}\frac{1}{n}\sum_{j=1}^{n}\left(\sqrt{|\boldsymbol{\alpha}_{j}^{*}\boldsymbol{z}|^{2}+{\varepsilon_{j}^{2}}}-\sqrt{b_{j}^{2}+\varepsilon_{j}^{2}}\right)^{2},$ (5.1)
where $\boldsymbol{\varepsilon}=[\varepsilon_{1},\cdots,\varepsilon_{n}]\in\mathbb{R}^{n}$ is a vector with prescribed values, subject to the condition that $\varepsilon_{j}\neq 0$ for all $b_{j}\neq 0$. When $\varepsilon_{j}=0$ for all $j=1,\cdots,n$, this model reduces to the complex amplitude-based model in [18], and is therefore referred to as the perturbed amplitude-based model. Based on this formulation, a new non-convex algorithm, called the perturbed Amplitude Flow (PAF) is developed. The PAF algorithm can recover the target signal under $\mathcal{O}(d)$ Gaussian random measurements. Starting from a well-designed initial point, PAF converges at a linear rate for both real and complex signals, without requiring truncation or reweighted procedures, while maintaining strong empirical performance.
@@
3:  Compute the initialization $\boldsymbol{z}_{0}=\lambda_{0}\cdot\boldsymbol{\nu}_{in}$..
@@ 6 Algorithm regarding to RGB picture
In quaternion image processing, the color channels (Red, Green, and Blue channels) are usually represented by the three imaginary components of quaternions, making the desired signal taking pure quaternions. Many studies (see e.g., [28]) opt to eliminate the real part of the quaternion signals after reconstruction. However, this approach generally performs poorly in the context of phase retrieval for pure quaternion signals. A more effective approach is to incorporate the pur quaternion assumption directly into the recovery process, as done in [15]. In this section, we brief review of the Phase Factor Estimate (QPFE) technique used in [15], which will be applied to refine our algorithms introduced in previous sections and then applied to phase retrieval of natural images.
@@
5:     Compute the left product of $\boldsymbol{z}$ by $\bar{q}$, i.e., $\boldsymbol{p}=\boldsymbol{z}\cdot\bar{q}.$
@@
8:  $\boldsymbol{z}_{T}$;
Similar to [15], we use $T_{p}$ to be the parameter that controls the transform of full quaternion to pure quaternion signals. Unlike [15] which utilizes inner and outer loop for the result, we only perform QPFE once after $T_{p}$ iteration periodically. In our algorithm, the total iteration number is the given parameter $T$ compare to the actual iteration number $T\times T_{p}$ in [15]. It is easily to see that the parameter $T_{p}$ makes difference when it changes, especially when $T_{p}=1$ which means operate QPFE in every iteration and $T_{p}=T$ which means operate QPFE at the final step of the full quaternion algorithm. Although there are difference among them, the actual computation shows no essential distinction. Since our work considers mainly on different algorithms rather than different aspects of a single algorithm, we prefer to save this part of work in the future.
@@ 7 Numerical Experiments
In this section, we show the numerical efficiency of QRAF and its improvements by comparing their the performance with other competitive methods. Throughout our work, the parameters of the tested algorithms are listed in Table 1. For the existing quaternionic algorithm QWF, we use the given parameters there. For the algorithms that transferred from the real or complex space, we compare the original parameters with a huge a mount of tested parameters to obtain a suitable set of parameters.
@@
Algorithm Parameters Algorithm Parameters
QWF $\eta=\frac{0.2n}{\sum_{k=1}^{n}y_{k}}$; QRAF $\left|{\mathcal{S}}\right|=\lfloor 3n/13\rfloor$, $\beta=5$, $\gamma=0.5$ and $\mu=6$
QRWF $\alpha_{\ell}=1$, $\alpha_{u}=5$ and $\mu=0.8$; QARAF $\eta=6$ and $\mu=0.8$;
QPAF $\gamma=1/2$, $\eta=2.5$ and $\sigma=2$; QAdRAF $\alpha=0.009$, $\mu=1$, $\epsilon=10^{-6}$
QTAF $\gamma=0.8$, $\eta=1.2$ and $\rho=\frac{1}{6}$ QIRAF $\eta=6$ and batch size $\left|{\Gamma}\right|=2^{k}$, $k=\lfloor\frac{\log\left(\frac{n}{4}-1\right)}{\log 2}\rfloor$;
@@ 7.1 Synthetic Data
i. Convergence performance   Firstly, we examine the convergence of each algorithm, as shown in Fig. 2 and Fig. 2. The relative error, defined as $\text{dist}(x,z)/\left\|{x}\right\|$ and $\text{dist}_{p}(x,z)/\left\|{x}\right\|$ w.r.t. full and pure quaternion algorithms respectively, serves as the convergence criterion. It is seen that all algorithms are convergent albeit at different rates. The convergence rate is a crucial aspect and is influenced not only by the algorithm’s gradient but also by other factors, such as the vector dimension $d$, the ratio $n/d$, and the algorithm’s initialization.
@@
We fix the oversampling rate to $n/d=9$ as an example to present the convergence analysis. It is easy to see that in general all algorithms converge to the same level of relative error, i.e. $\ln(10^{-35})$, but there are some details that must be noted. For the full quaternion algorithms in Fig. 2, QARAF leads the competition of convergence with the iteration number less than $400$. QRAF and its variants converge evidently faster than other algorithms. The QWF also converges but with a lower speed that is not fully shown in the figure. We should note that the curves of QRAF and QIRAF almost cover each other, showing the same ability of convergence. This is actually important since as the variants of QRAF, QIRAF has the advantage of computation speed which we will discuss later in details.
@@
ii. Success rate  We next evaluate the success rates of QRAF and its variants, along with other benchmark algorithms, under varying $n/d$ rate. The oversampling rate $n/d$ is a critical factor in phase retrieval, as lower values imply reduced computational cost, faster execution, and lower memory consumption. Therefore, achieving successful recovery at smaller $n/d$ is a key indicator of algorithmic efficiency.
We conduct experiments on quaternion-valued signals $\boldsymbol{x}\in\mathbb{H}^{d}$ with $d=64$ and $d=100$, respectively. For each configuration, we perform 100 trials over $n/d\in[[3,13]]$ with increments of $0.5$. Noting that the interval $n/d\in(5,8)$ is particularly sensitive, we reduce the step size to $0.2$ within this range to capture more granular performance behavior. Each algorithm was run for 1500 iterations per trial. Similar as in [15, 14], a trial was deemed successful if it achieved $\text{dist}(\boldsymbol{x},\boldsymbol{z})<10^{-5}$ within the 1500 iterations. The hyperparameter settings for each algorithm are listed in Table 1, and the resulting success rates are reported in Fig. 4 and 4. The most notable performance is achieved by QARAF, which reaches a success rate of $1.0$ with an oversampling ratio below $n/d=7$, clearly outperforming all other algorithms. This superior performance is attributed to the accelerated steepest descent strategy adopted in QARAF, as detailed in Section 4.2.
Figure 3: Success rate for $d=64$
Figure 4: Success rate for $d=100$
QAdRAF, which incorporates an adaptive gradient descent mechanism, also demonstrates improved performance over both the baseline QRAF and other methods in the benchmark group shown in Fig. 4. Particularly noteworthy is the performance of QIRAF. While its success rate closely matches that of QRAF, it substantially reduces computational overhead by employing a mini-batch strategy. Despite this simplification, QIRAF retains nearly the same recovery capability, making it an attractive alternative for practical implementations. As shown in Table 2, QIRAF achieves a comparable runtime to QARAF, thus offering an effective trade-off between accuracy and efficiency.
Two important remarks apply to the interpretation of success rate curves. First, since success is defined solely by reaching $\text{dist}(\boldsymbol{x},\boldsymbol{z})<10^{-5}$ within 1500 iterations, the curves do not distinguish between fast and slow convergence within a trial. Second, small fluctuations or inflection points in the curves (e.g., due to randomness in measurement matrices) are expected, especially when $n/d$ varies in small increments (e.g., 0.2). These do not affect the overall trends and conclusions, and increasing the number of trials or step sizes would smooth the curves further. Therefore, the current setting is considered sufficient for comparative evaluation.
@@
From Fig. 5, we can easily seen that the signal dimension also affect tested algorithms on convergence at some level. For instance, when $d=30$ and $n/d=6$, the success rate of QIRAF is $0.83$ while when $d=300$ the success rate is 00 with the same $n/d$. Another example is when $d=60$ and $n/d=6.5$ the success rate is $1$, but when $d=300$ and $n/d=6.5$, the success rate is $0.33$. (The original data is uploaded and can be verified.) In another word, as $d$ increases, the ratio $n/d$ must also be larger to achieve a high success rate. It is worth noting that, compared to QRAF, the QIRAF is more sensible to the size of vector $\boldsymbol{x}\in\mathbb{H}^{d}$. And the QARAF is the most stable one, comparing to its counterparts.
@@ 7.2 Experiment: Color Images
In this section, we compare different quaternionic algorithms in processing real images. The images used for testing are sourced from The USC-SIPI Image Database  [^1] https://sipi.usc.edu/database/, i.e. 4.1.04 ($256\times 256$), 4.1.05 ($512\times 512$) and 4.2.03 ($256\times 256$). Ideally, the entire image would be considered as the signal for reconstruction. However, due to computational limitations present not only in quaternionic algorithms but also in real and complex algorithms, we adopt the conventional approach of segmenting each image into smaller blocks and treating each block as a separate signal $\boldsymbol{x}$ to be reconstructed. Specifically, in the first experiment, i.e. Fig. 6, we test the picture 4.1.04 in the database as in [15]. The picture is $256\times 256$, we split this picture into $32\times 32$ pieces, then each piece is a $8\times 8$ block containing $64$ pixel which is the signal $\boldsymbol{x}$ and its size $d=64$. In the second experiment, i.e. Fig. 7, the figure 4.1.05 of the size $512\times 512$ is also tested. The figure is divided also into $32\times 32$ pieces but with each block of $16\times 16$ pixels. Therefore, in this experiment, the size of the signal $\boldsymbol{x}$ is $d=256$.
For these experiments, we set the $n/d$ ratio to $9$ and fixed the total number of iterations to $T=300$ for each algorithm. The outcomes are presented in Fig. 6. The results in Fig. 6 clearly highlight the performance of each algorithm. In comparison with the original image, only the reconstruction by PQRAF (see Fig. 6f) contains no defective blocks, indicating the robustness of this algorithm. PQPAF and PQRWF also perform well, with only a few defective blocks, reflecting the varying strengths of WF and AF. Conversely, PQWF and PQTAF exhibit relatively weaker performance, illustrating that different WF and AF algorithms have trade-offs in terms of convergence speed and overall effectiveness.
@@
To quantify these observations, we utilize two evaluation metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). PSNR measures the similarity between the original and reconstructed images, with higher values indicating closer similarity. SSIM assesses structural similarity, ranging from 00 to $1$, where a score of $1$ denotes perfect similarity. The indices reveal that PQRAF significantly outperforms other algorithms, achieving the highest PSNR score of $96$ in this experiment. Additionally, both PQRAF and PQPAF attain the maximum SSIM score of $1$, leading the competition in this aspect. These indices corroborate the earlier qualitative observations with a more precise, quantitative evaluation.
@@ 7.3 Comparison of conventional and quaternionic algorithms
For a pure quaternion signal $\boldsymbol{p}\in\mathbb{H}^{d}_{p}$, the RAF of a monochromatic model separately reconstruct $\mathcal{P}^{i}(\boldsymbol{p})$, $\mathcal{P}^{j}(\boldsymbol{p})$ and $\mathcal{P}^{k}(\boldsymbol{p})$, which are three real signal in $\mathbb{R}^{d}$. We denote by $\hat{\boldsymbol{z}}^{i}$, $\hat{\boldsymbol{z}}^{j}$ and $\hat{\boldsymbol{z}}^{k}$ the reconstructed signal corresponding to each channel. To keep the metric consistent, the reconstruction error for real RAF of a quaternion signal $\boldsymbol{p}$ is measured as $\left(\sum_{h=i,j,k}{\rm dist}_{p}\left(\mathcal{P}^{h}(\boldsymbol{p}),\hat{\boldsymbol{z}}^{h}\right)^{2}\right)^{1/2}$. On the other hand, the concatenation model of real RAF for a pure quaternion signal $\boldsymbol{p}\in\mathbb{H}^{d}_{p}$ also split the signal as $\mathcal{P}^{i}(\boldsymbol{p})$, $\mathcal{P}^{j}(\boldsymbol{p})$ and $\mathcal{P}^{k}(\boldsymbol{p})$, and concatenate the three real vector into a new signal, i.e. $\hat{\boldsymbol{p}}=\left[\left(\mathcal{P}^{i}(\boldsymbol{p})\right)^{T},\left(\mathcal{P}^{i}(\boldsymbol{p})\right)^{T},\left(\mathcal{P}^{i}(\boldsymbol{p})\right)^{T}\right]^{T}\in\mathbb{R}^{3d}$. We denote by $\hat{\boldsymbol{z}}\in\mathbb{R}^{3d}$ the reconstructed signal, then the measurement for reconstruction error of this algorithm is defined as ${\rm dist}_{p}\left(\hat{\boldsymbol{p}},\hat{\boldsymbol{z}}\right)$.
@@
However, one thing that cannot be neglected is that real RAF algorithms are evidently faster than quaternionic ones in reconstructing an image as we already mentioned. By simply increasing the $n/d$ rate without any time consumption piles up, real algorithms can surpass the quaternionic ones in convergence as well. However, such comparisons would no longer adhere to the same standard, and thus fall outside the scope of this study.
8 Conclusions
In this work, we addressed the quaternionic phase retrieval (QPR) problem by proposing the Quaternionic Reweighted Amplitude Flow (QRAF) algorithm, a novel generalization of the real/complex amplitude flow framework to the quaternion domain. Our method innovatively extends the reweighted gradient descent paradigm by carefully accounting for the noncommutative nature of quaternion multiplication. In addition, we developed three algorithmic variants of QRAF to further enhance convergence and stability. Comprehensive experiments on both synthetic datasets and real-world color images demonstrate that QRAF consistently outperforms existing quaternionic methods such as QWF, QTWF, and QTAF in terms of convergence speed, recovery accuracy, and robustness. In particular, our methods achieve higher success rates under limited measurements, and demonstrate superior PSNR and SSIM scores in practical image reconstruction tasks.
To further improve theoretical rigor, we introduced the Quaternionic Perturbed Amplitude Flow (QPAF) algorithm, which is based on a perturbed amplitude model and inherits provable linear convergence properties from its complex counterpart. While QPAF achieves slightly lower empirical performance than QRAF, its simplicity and theoretical guarantees make it a valuable complementary approach.
9 Future Work
While the proposed methods achieve strong empirical results, several questions remain open for future investigation.
(1) Theoretical analysis of QRAF: One notable challenge is the absence of a proof for the Local Convergence Radius (LCR) condition in the QRAF, as outlined in Problem 3.5. This limitation is not unique to the quaternionic setting but is also observed in the complex setting. We anticipate that future work establishing the LCR condition for complex signals will facilitate its extension to the quaternionic framework without significant complexity.
(2) Efficient quaternionic computation: The computational cost of QPR algorithms remains high compared to their real-valued analogues. This is largely due to the lack of hardware-accelerated support and optimization frameworks for quaternion algebra. Future work may explore dedicated quaternionic linear algebra libraries, GPU-accelerated quaternion solvers, or fast approximation schemes that retain algebraic structure while reducing computational complexity.
(3) Parameter tuning strategies: The performance of QRAF and QPAF is sensitive to initialization and step-size parameters, which are less well understood in quaternionic optimization. Developing adaptive parameter selection schemes—such as learning-based hyperparameter tuning or curvature-aware step size adjustment—could improve robustness across diverse signal types. It is also an interesting topic.